OpenStudy (anonymous):

y=(x−5)2−4 Identify the vertex, equation for the axis of symmetry, domain, range and x- and y-intercepts of the parabola. y=−2(x+8)2+2 Identify the vertex, equation for the axis of symmetry, domain, range and x- and y-intercepts of the parabola.

OpenStudy (anonymous):

y=(x−5)2−4 y+4 = (x−5)2 vertex (5,-4)

OpenStudy (anonymous):

vertex = -b/2a domain is = x range = y x intercept is when y is = 0 y intercept is when x is = 0